OpenStudy (anonymous):

Please help me to simplify the expression without using negative exponents. (x^1/2 y^5)^-4/3

OpenStudy (anonymous):

\[(x^{1/2} y^{5})^{-4/3}\] By raising the term inside the parenthesis to the -4/3 power, we are multiplying both exponents by -4/3. So we'll get: \[x^{1/2 \times -4/3} y^{5 \times -4/3}\] or: \[x^{-4/6} y^{-20/3} = x^{-2/3}y^{-20/3}\] If you want to eliminate the negative exponents, move both terms into the denominator: \[\frac{ 1 }{ x^{2/3} y^{20/3} }\]

OpenStudy (anonymous):

I got x^5/3/y^1/3 did I simplify it correctly?

OpenStudy (anonymous):

Just saw your answer and I was way off. Thank you for the steps.